OpenStudy (anonymous):

You are decorating a cylindrical hat box to give to your grandmother as a gift. You want to paste family photos onto the curved part of the box. The height of the box is 8 inches and the radius is 6 inches. If none of the photos overlap, what is the area of photographs you need? Use 3.14 for pi .

OpenStudy (ghazi):

2*(pi*r*l)= 2*(3.14*6*8)

OpenStudy (ghazi):

also if you're including the faces of cylinder then you'll have to add 2*pi*r^2

OpenStudy (ghazi):

and net surface area= 2*3.14*(6)^2+ 2*3.14*6*8

OpenStudy (anonymous):

301.44 is that the awnser ?

OpenStudy (anonymous):

301.44 in sq